10. The curve y = cot x, the line x = 4, and the x-axis. (A) In 2 (B) 2In1 (C) 1 (D) In 2 (E) 2
Answer #1

The given region can be plotted as

1 5 0 5 2 5 1 5 0

The area of the region is given by

π/4 cot x.dr

integrate wrt x

A = [in |sin·rliğ/4

A = \ln \left | \sin \frac{\pi}{4}\right |

4 - In y2 V/2

A--In V2

ln 2

Which might be wrong.

But please let me know if I have to find the volume of rotation here.

You didn't mention what to find here.

I hope this answer helps,
Thanks,
Keep posting questions.